OpenStudy (candlelights):

How should someone start trying to find the solutions of sin 2 x - sin 4 x? I have some ideas on how to start, but I'm not sure what would be the right way to approach this. Would you extract sin 2 x to get sin 2 x (sin x - sin 2 x) and go from there, or would you simplify sin 2 x to 2 sin x cos x to get (2 sin x cos x) - sin 4 x? And could you explain why that's so?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\sin^2(x)-\sin^4(x)\]?

OpenStudy (owen3):

sin(4x) = sin(2*2x) = 2 sin(2x) cos (2x)

satellite73 (satellite73):

not really clear, and in any case it is not an equation, so there is nothing to "solve"

OpenStudy (misty1212):

maybe \[\sin^2(x)-\sin^4(x)=0\]? not clear you got a screen shot?

OpenStudy (candlelights):

@misty1212 Nope. It'd be \[\sin 2 x - \sin 4 x\] @satellite73 what do you mean it can't be solved? It does have multiple solutions, so there's no one solution if that's what you meant. I'm working within the intervals [0, 2pi] if that helps.

satellite73 (satellite73):

you can solve an equation like \[2x-1=0\] you cannot solve \[2x-1\]

OpenStudy (candlelights):

@satellite73 oh, i get what you mean. yeah. in my case sin 2 x - sin 4 x does equate to 0.

OpenStudy (mathmale):

The problem is, your mathematical statement is NOT an equation and thus cannot be solved. If you append "=0" to your statement, then it becomes solvable.

OpenStudy (mathmale):

Putting that "=0" in there is NOT optional. It's mandatory.

OpenStudy (candlelights):

@mathmale I see. I should've worded that properly, my bad.

OpenStudy (mathmale):

Glad to hear this.

OpenStudy (mathmale):

Note that sin 4x can be re-written as sin (2x + 2x). Look up the summation formula for the sine function: sin 2x = ???

OpenStudy (mathmale):

Then apply that formula to sin 4x.

OpenStudy (candlelights):

@mathmale Summation formula? I don't think my school mentioned anything like that! Well, I can work with this. This makes my equation \[\sin 2 x - \sin (2x + 2x)\], right? So from here, do I redistribute the sin 2 x to get something like:\[\sin 2 x - \sin 2 x + \sin 2 x\]?

OpenStudy (mathmale):

Look up the formula for sin (a + b). That's the "summation formula" I'm referring to.

OpenStudy (candlelights):

Wait a minute, it'd be \[\sin 2 x - \sin 2 x - \sin 2 x\] because of the -1

OpenStudy (mathmale):

Not so. Please look up "trig formulas" and don't risk getting led astray by guessing.

OpenStudy (candlelights):

OK, I see. Would you use a sum and difference formula here?

OpenStudy (mathmale):

sin (a+b) = sin a cos b + cos a sin b. You'll need this formula in the future; may as well make a note of it and review it every now and then. Yes, you could call this the "sum and difference formulas" for the sine.

OpenStudy (mathmale):

Apply this formula to expand sin (4x) = sin (2x + 2x).

OpenStudy (candlelights):

Ah, so they did teach this to me! we simply called a and b x and y. OK, This gives me: \[\sin 2 x - \sin 2 x \cos 2 x + \cos 2 x \sin 2 x\]

OpenStudy (mathmale):

Have to put parentheses around the last 2 terms:\[\sin 2 x - [\sin 2 x \cos 2 x + \cos 2 x \sin 2 x]\]

OpenStudy (mathmale):

Note that sin 2x is now a factor of all three terms. What next?

OpenStudy (candlelights):

Factor out sin 2 x?

OpenStudy (mathmale):

Be careful with your signs. Those parentheses do make a difference. Yes. Factor sin 2x out. and simplify the results.

OpenStudy (candlelights):

I know this sounds silly, but can I ask how you'd factor out this kind of equation? I haven't tried to factor out something like this before, I believe. Would it go something like this?: sin 2 x (-[cos 2 x + cos 2 x]) This is just my guess, but of course I can't be sure that's right.

OpenStudy (mathmale):

sin 2x divided by sin 2x is 1, right? Where's your '1'? What is -(cos 2x + cos 2x)?

OpenStudy (candlelights):

Oh, OK. So should this be \[\sin 2 x (1 - (1 \cos 2 x + 1 \cos 2 x)\] ?

OpenStudy (mathmale):

Yes, that's partially right. But what is 1cos 2x + 1cos 2x?

OpenStudy (candlelights):

2 cos 4 x?

OpenStudy (mathmale):

What is 1 egg + 1 egg? Same principle applies to 1cos 2x + 1cos 2x.

OpenStudy (candlelights):

So, 2 cos 2 x.

OpenStudy (candlelights):

And the cos 2 x doesn't change when you add 2 of the same kind together? Of course.

OpenStudy (mathmale):

Much better. Then you have sin 2x (1-2cos 2x). You must set this =0 and then solve for x. There will be more than one solution.

OpenStudy (candlelights):

Yup, this was what I needed. Thanks!

OpenStudy (mathmale):

My pleasure. Good luck.